By taking an airplane Larry can travel from Denver to Houston in one hour. The same trip takes 5 hours by bus. Airfare is $90 and the bus fare is $30.
Larry, when he is not traveling, can work and earn $30/hour
1
Expert's answer
2013-02-27T06:25:28-0500
OC = (Direct Cost) + (Indirect Cost) = (Bus Fare) + (Time Cost)

= $30 + (5 hours) ($30/hour)

= $30 + $150 = $180

OC = $90 + (1 hour)($30) = $120

Plane is cheaper for Larry

For Moe,

OC (bus) = $30 + (5 hours)($6/hour) = $60

OC (plane) = $90 + (1 hour) ($6/hour) = $96

Therefore, bus is cheaper for Moe.

Lesson:

People with high time value (e.g. busy executives) are more likely to travel by
plane than by bus; whereas, people with low time value (seniors, students on
vacation) are more likely to travel by bus.